DIRECTIONS: In each problem below, give the null and alternative hypothesis and

identify whether it is right-tailed, left-tailed or two-tailed test.

3. A quality control engineer is testing the battery life of a new smartphone. The company

is advertising that the battery lasts 24 hours on full- charge, but the engineer suspects that the

battery life is actually less than that. They take a random sample of 50 of these if their average

battery life is significantly less than 34 hours.

4. In the past, the mean running time for a certain type of radio battery has been 9.6 hours.

The manufacturer has introduced a change in then production method and wants to perform a

hypothesis test to determine whether the mean running time has changed as a result.

5. In a random sample of 400 electronic gadgets, 14 were found to be defective. The

manufacturer wants to claim that more than 5% of all the gadgets are defective. Test this claim at

the 0.01 level of significance.


1
Expert's answer
2022-05-05T17:11:18-0400

3. The null hypothesis states that the battery life of a new smartphone will be greater than or equal to 24 hours.

The alternative hypothesis states that the battery life of a new smartphone will be less than 24 hours.

"H_0: \mu\ge24"

"H_1: \mu<24"


B. This corresponds to a left-tailed (directional, one-tailed) test.


4. The null hypothesis states that the mean running time for a certain type of radio battery will be equal to 9.6 hours.

The alternative hypothesis states that the mean running time for a certain type of radio battery will not be equal to 9.6 hours.

"H_0: \mu=9.6"

"H_1: \mu\not=9.6"


B. This corresponds to a two-tailed test.


5. The following null and alternative hypotheses need to be tested:

"H_0:p\le0.05"

"H_a:p>0.05"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z: z > 2.3263\}"

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.035-0.05}{\sqrt{\dfrac{0.05(1-0.05)}{400}}}=-1.376"

Since it is observed that "z=-1.376<2.3263=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, is "p =P(Z>-1.376)=0.915589," and since "p= 0.915589>0.01=\alpha," it is concluded that the null hypothesis is notrejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is more than 5%, at the "\alpha = 0.01" significance level.